OpenStudy (anonymous):

Using complete sentences, explain the key features of the graph of the tangent function.

OpenStudy (anonymous):

I gave them to you on the last post.

OpenStudy (anonymous):

Look at the properties of tan

OpenStudy (anonymous):

that didnt help lol

OpenStudy (anonymous):

Sure it did. Those are all the key features.

OpenStudy (anonymous):

can you say them again?

OpenStudy (anonymous):

If you can name one I will give you the list.

OpenStudy (anonymous):

asymptote

OpenStudy (anonymous):

Right!!! Here are the properties of the tan The domain is the set of all real numbers except odd multiples of \( \huge \frac{\pi}{2} \) The range is the set of all real numbers The tangent function is an odd function, as the symmetry of the graph with respect to the origin indicates. The tangent function is periodic with period pi The x-intercepts are .....,-2pi, -pi, 0, pi, 2pi, 3pi, ..... Vertical asymptotes occur at x = odd multiples of \( \huge \frac{\pi}{2} \)

OpenStudy (anonymous):

Here's the answer I put The key features of the graph of the tangent function are the starting point, asymptotes and the intercepts. These features are the main parts of the function that connect the sine and cosine.

OpenStudy (anonymous):

Everything on that list are the key features

OpenStudy (anonymous):

You need to know all that when it comes to tan and graphing it.